OpenStudy (anonymous):

The area of a circle with radius R is given by the formula A=πr².If a circle has area 225π square inches, what is the radius?

OpenStudy (katrinakaif):

A = pie r^2 225 pie = pie r^2 (Pie and pie cancel each other out when divided leaving us with..) 225 = r^2 and take the square root 15 = r

OpenStudy (anonymous):

\[πr^2 = π225r^2 ⇒ r^2 = 225\] Then as KatrinaKaif says, take the square root of both sides, _et voilà_.

OpenStudy (anonymous):

Course, I made a mistake above… Wish I could edit… but it should not say \[\pi225r^2\] but rather \[\pi225\]

OpenStudy (anonymous):

inother words the answer is 15sqrtpie

OpenStudy (anonymous):

No, just 15, @outkast3r09. The question was what is the radius, not what is the square root of the area.